OpenStudy (anonymous):

Find all the roots of the equation \(x^5-1=0\).

OpenStudy (phi):

write 1^(1/5) as e^(i*2pi)/5 then take move around the unit circle to get the rest of the roots.

OpenStudy (mani_jha):

Is the answer 1?(only one root)

OpenStudy (anonymous):

1 real root, 5 complex ones

OpenStudy (anonymous):

5 or 4?

OpenStudy (anonymous):

5, including the real one

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Isn't that the fundamental theorem of algebra?

OpenStudy (anonymous):

idk what fundamental theorem of algebra is?

OpenStudy (experimentx):

I though there was some kind of factorization rule for x^n - a^n = (x-a)(....) ... nevertheless, the method by phi is far superior.

OpenStudy (mani_jha):

Oh, that was obvious. Degree is 5, so a total of 5 roots.

OpenStudy (anonymous):

Can you find the other roots? (2pi/5+2pi), ...?

OpenStudy (anonymous):

wolframalpha style? http://www.wolframalpha.com/input/?i=x%5E5-1%3D0

OpenStudy (anonymous):

Yeah, I could have done that. :/ silly me. Thanks btw.

OpenStudy (experimentx):

interesting.

OpenStudy (anonymous):

.....

OpenStudy (phi):

in case it is not clear, you move in increments of 2pi/5 : you get (as the exponent to e) 2pi/5 , 4pi/5 , 6pi/5, 8pi/5, 10pi/5 the last is just 2*pi and you are at 1 on the real axis

OpenStudy (anonymous):

\[\large x^5-1 = 0\] \[\large \implies (x - e^{\frac{i2\pi}{5}})(x - e^{\frac{i4\pi}{5}})(x - e^{\frac{i6\pi}{5}})(x - e^{\frac{i8\pi}{5}})(x - 1)=0\] So, can I do this? \[\int \frac{1}{x^5-1}dx = \frac{1}{(x - e^{\frac{i2\pi}{5}})(x - e^{\frac{i4\pi}{5}})(x - e^{\frac{i6\pi}{5}})(x - e^{\frac{i8\pi}{5}})(x - 1)}dx\]

OpenStudy (phi):

if x^5-1 = 0 isn't this undefined?

OpenStudy (anonymous):

The problem is to integrate the function \(\large\frac{1}{x^5-1}dx\). \[x^5-1 = (x - e^{\frac{i2\pi}{5}})(x - e^{\frac{i4\pi}{5}})(x - e^{\frac{i6\pi}{5}})(x - e^{\frac{i8\pi}{5}})(x - 1)\]Maybe I can use this to integrate it.

OpenStudy (anonymous):

Partial fractions. but I'm not sure.

OpenStudy (anonymous):

I haven't tried anything like this before.

OpenStudy (phi):

It looks like a lot of work. post your answer if you can get to one.

OpenStudy (anonymous):

The answer on wolframalpha is pretty huge D: